OpenStudy (anonymous):

x + 40 divided by 4 = 45 degrees

OpenStudy (anonymous):

35

OpenStudy (anonymous):

x+40/4=45 You divide 40/4= 10 x+10=45 is your new equation Minus 10 from each side x+10-10=45-10 x=35